2. Monster Inc., the maker of personal computer games has been contacted by Hot Stuff Inc., which has developed a new technology for personal computer gamers which provides a new challenge for gamers. The new technology consists of a seat pad for the gamer’s chair which is connected to a control unit which is connected to the personal computer. When the gamer makes an error while playing Daytona 500 he/she receives a “mild” electric shock through the seat pad. The new technology, which has received extremely positive reviews from a panel of gamers, consists of the Daytona 500 game CD, the seat pad, the control unit and cables. In tests, by Monster Inc., the new technology has performed flawlessly, is considered safe and Monster Inc. is considering licensing Hot Stuff’s new gamer technology. The following cost information pertains to the new technology: Hot Stuff Daytona 500 Software License $200,000 New Equipment To Produce Seat Pads $50,000 Advertising and promotion $150,000 Material & Royalties “Daytona 500” Game (per 100 CDs) $500 Material & Labor For Seat Pad (per 10 pads) $150 Control Unit And Cables (per unit) $5 Market research estimates that the market for the new Daytona 500 game and “Hot Stuff” is 1,000,000 units at a retail price of $75, which includes the game CD, seat pad, control unit and cables. Monster Inc. traditionally sells through major retailers like CompUSA. The retailer’s traditional margin is 20%.

a. How much revenue does Monster Inc. receive for each unit sold?

b. What is Monster Inc.’s contribution margin per unit?

c. What is the break-even volume in units?_____________________________

d. How many units did Monster sell if Total Sales Revenue = $1,000,000?

e. What is the profit if Monster’s Total Sales Revenue = $1,000,000?
Answer #1
a) Unit revenue to Monster Inc. = selling price * [1 - ( 100 / (100 + traditional margin))]
75 - (75*100 /120) = $62.5
b)
Variable Cost per unit =
CD cost = $500 / 100 CD = $5 per CD
Seat Pad = $150 / 10 pads = $15 per pad
Control unit & cable = $5
Total Variable cost per unit = $25
Contribution per unit to Monster Inc. =
Unit revenue - Variable cost per unit = $62.50 - $25 - $37.5
c)
Fixed Costs = $200000 + $50000 + $150000 = $400000
Break even sales (in units) = FC / Contribution per unit
400000 / 37.5 = 10667 units
d) Units sold for total revenue $1m = total revenue / Unit revenue
1000000 / 62.5 = 16000 units
e) Profit if Monster's Total revenue $1m =
1000000 * (37.5/62.5) - 400000 = 200000